OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (anonymous):

\[|x+2| = -2\] the statement is false

OpenStudy (jamesj):

@newmonei : you can get people to give you the answers to all these questions. But if you don't understand the concept here which we're trying to ask you about, then you will not understand those answers and it's a pointless exercise.

OpenStudy (anonymous):

|x+2|=14-16=-2 removing modulus (x+2)=-2 or -(x+2)= -2 x=-2-2=-4 or (x+2)=2,x=0 x=-4,0 ans

OpenStudy (jamesj):

@mukesh : you don't understand it either because your answer is completely wrong.

myininaya (myininaya):

can distance be negative?

OpenStudy (jamesj):

ok, newmonei..., you're not responding, I'm out of here. Good luck to you.

OpenStudy (anonymous):

Negative distance... interesting idea.

OpenStudy (anonymous):

.....